OpenStudy (salazarblack):

number of solutions for: sin^2+cosx+1=0 in interval (0,4pi)

OpenStudy (salazarblack):

@iPwnBunnies

OpenStudy (salazarblack):

i have a problem 'cause even when i know how to simplify the whole thing i don't get how do i get to the solutions such as x= pi/2 +kpi or something like that :|

OpenStudy (salazarblack):

@ganeshie8

OpenStudy (mathmale):

There is an identity for sin^2 (which should be written as (sin x)^2, by the way) that would enable you to re-write the given equation in terms of the cosine function alone. What is that identity? please apply it here.

OpenStudy (salazarblack):

i need to tell you that i'm completely clueless when it comes to trigonometry >_<

OpenStudy (salazarblack):

should i add 2sinxcosx so that i can make it as (sinx +cosx)^2 or? o.O

OpenStudy (mathmale):

I'm sorry you feel "clueless" when it comes to Trig, Salazar, but I most certainly expect that you will do something to help yourself in solving this problem. I hinted that there is an identity for (sin x)^2 and ask that you find and present that identity here.

OpenStudy (mathmale):

Have you access to a table or list of trig identities? If not, I'd strong ly sugges you begin making up one now for continual reference.

OpenStudy (salazarblack):

i do but all i can think of using here is 1-cos^2x but then i get 2=0 which makes no sense..

OpenStudy (dls):

sin^2x+cosx+1=0 (1-cos^2x)+cosx+1=0 cosx-cos^2x=0 cosx(1-cosx)=0 cosx = 0 or cosx = 1 x = pi/2 or x=0 now see in interval (0,4pi)

OpenStudy (salazarblack):

ohh right i have cosx not cos^2x so they don't shorten

OpenStudy (mathmale):

The trig identity relating (sin x)^2 and (cos x)^2 is ... what? You fill in the blanks, please. It's important that you represent the square of the cosine as either (cos x)^2 or \[\cos^2 x\]

OpenStudy (mathmale):

Salazar: Stop a moment. Re-read what DLS and I have typed to you. What further info need you to solve this problem yourself? Be specific, please.

OpenStudy (salazarblack):

1? and all i need is a number of solution that work for the (0,4pi ) interval

OpenStudy (salazarblack):

solutions*

OpenStudy (mathmale):

I see you may be experiencing an overload of help here. You're starting with the equation (sin x)^2+cosx+1=0 and must use the identity I've been discussing to change that (sin x)^ into an equivalent function in cos x alone. What is that identity? It's essential that you know it and be able to apply it. Using this identity, re-write (sin x)^2+cosx+1=0 in terms of cos x alone.

OpenStudy (salazarblack):

well isn't that 1-cos^2x? o.O

OpenStudy (mathmale):

Please type that as 1 - (cos x)^2 for clarity. Yes. Use this identity to write the original equation in terms of cos x only, please.

OpenStudy (salazarblack):

1-(cos x)^2 +cos x +1=0 and i did understand what DLS wrote

OpenStudy (mathmale):

All right. simplify 1-(cos x)^2 +cos x +1=0 now, please.

OpenStudy (salazarblack):

cos x -(cos x)^2 +2 =0

OpenStudy (mathmale):

Good. that can be simplif. further to read (cos x)^2 - cos x - 2 = 0. Verify that this is correct or explain where you think it's wrong.

OpenStudy (salazarblack):

it's correct

OpenStudy (mathmale):

Are you able to factor that? Hint: you might want to factor x^2 - x - 2 = 0 first, for practice.

OpenStudy (salazarblack):

(x-2)(x+1)

OpenStudy (mathmale):

(x-2)(x+1)=. Therefore, x = ? and x=?

OpenStudy (salazarblack):

x=2 or x=-1

OpenStudy (mathmale):

Good. Now replace that x with "cos x". cos x = ? and cos x = ?

OpenStudy (salazarblack):

cos x =2 and cos x=-1

OpenStudy (mathmale):

Are there any angles between 0 and 4 pi at which cos x = -1? If so, at which x? (x represents an angle).

OpenStudy (mathmale):

If you're not sure, you could always draw or look up the graph of y = cos x and then determine where cos x = -1.

OpenStudy (salazarblack):

-pi .. -180

OpenStudy (salazarblack):

it should be -pi +2npi but i don't get how do i know it's 2npi?

OpenStudy (mathmale):

I'd prefer you state, "cos x = -1 in only one place on the interval [0, 2pi] (or [0,360]" cos x=-1 .... where? ... on the interval [0, 4pi]

OpenStudy (salazarblack):

720? o.O

OpenStudy (mathmale):

Is that correct? is cos 720 = -1?

OpenStudy (mathmale):

Again, I'd suggest you obtain the answers needed by graphing cos x on [0,4pi].

OpenStudy (mathmale):

is cos 720 = -1? no.

OpenStudy (salazarblack):

i guess not :D

OpenStudy (mathmale):

so, salazar, cos x = -1 at x = pi radians and then again at x = ???

OpenStudy (mathmale):

This is an essential skill to have, so I'm going to stick to this until you come up with a solution to cos x = -1 in the interval [2pi, 4pi].

OpenStudy (salazarblack):

let me draw that

OpenStudy (mathmale):

Just draw the graph of the cosine function on [0,2pi] and then extend it to 4pi. This will give you TWO full cycles of the cosine function. Again, this is essential to know.

OpenStudy (salazarblack):

well the line goes through 2pi

OpenStudy (mathmale):

Salazar, as we've already seen, the solution (only solution) of cos x = - 1 on [0,pi] is pi. The cosine function is periodic with period 2pi, and thus, if we add 2pi to pi, we get 3 pi. Thus, the roots of cos x = -1 on [0,4pi] are ... { , } (fill in the blanks)

OpenStudy (salazarblack):

(0,5pi)

OpenStudy (mathmale):

No, afraid not; the roots are pi and 3pi: {pi, 3pi}.

OpenStudy (mathmale):

Likewise, if you extend the interval again , to [0,6pi], the roots as {pi, 3pi, 5pi}. How far apart are those roots?

OpenStudy (salazarblack):

what do you mean "how far apart"? how far in pi or..?

OpenStudy (mathmale):

3p-1pi=? 5pi-3pi=?

OpenStudy (salazarblack):

oh 2 and 2

OpenStudy (mathmale):

I appreciate your asking good questions here. Actually, not 2 and 2, but 2pi and 2pi, OK?

OpenStudy (salazarblack):

okay

OpenStudy (mathmale):

We've found several roots, right? the 1st positive root is x = pi. Starting with x = pi, add 2pi, then add 4pi , then add 6 pi. what do you get? pi, 3pi, ..., ...,

OpenStudy (mathmale):

Sal? pi, 3pi, ? , ? , ....

OpenStudy (mathmale):

Sorry to make you repeat this over and over, but I want you to see the pattern that is developing.

OpenStudy (salazarblack):

okay pi,3pi,7pi,13pi

OpenStudy (mathmale):

Wouldn't that be pi, 3pi, 5pi, 7pi, 9pi, ...?

OpenStudy (mathmale):

Why 7pi, 13pi, with no 9pi, no 11pi?

OpenStudy (salazarblack):

you said ad 4 and then 6...

OpenStudy (salazarblack):

add*

OpenStudy (mathmale):

But anyway. So, if you choose x = -1 as your beginning root, the next root is x=-1+2pi; the next root is -1+4pi, the next is -1+6pi, and the next is .... ??

OpenStudy (salazarblack):

-1 + 8pi?

OpenStudy (mathmale):

Exactly. And therefore, your general solution is x = -1 (plus or minus)n*2pi, where n=0, plus or minus 1, plus or minus 2, and so on. If you want the solutions ONLY on [0,4pi], let n=0 and 1 and you will end up with pi+0*2pi = pi pi+1*2pi = pi + 2pi = 3pi (which we already knew).

OpenStudy (mathmale):

Does this conversation answer your questions / concerns?

OpenStudy (salazarblack):

okay okay , got it xD thank you :)

OpenStudy (mathmale):

My pleasure. Thanks for sticking with me and not complaining ab out my going on and on. I truly wanted you to undrstand the concepts behind this problem.

OpenStudy (salazarblack):

np :D